### v1

# HARBOR: Holistic Adaptive Risk assessment model for BehaviORal healthcare

Abstract
Behavioral healthcare risk assessment remains a challenging problem due to the highly multimodal nature of patient data and the temporal dynamics of mood and affective disorders. While large language models (LLMs) have demonstrated impressive reasoning capabilities, their effectiveness in structured clinical risk scoring remains unclear. In this work, we introduce HARBOR, a Behavioral Health–aware language model designed to predict a discrete mood and risk score, termed the Harbor Risk Score (HRS), on a Likert scale from 3-3 (severe depression) to +3+3 (mania). We also release PEARL, a longitudinal behavioral healthcare dataset spanning four years of monthly observations from three patients, containing physiological, behavioral, and self-reported mental health signals. We benchmark traditional machine learning models, proprietary LLMs, and HARBOR across multiple evaluation settings and ablations. Our results show that HARBOR substantially outperforms both classical baselines and off-the-shelf LLMs, achieving a 69% accuracy compared to 54% for logistic regression and 29% for the strongest proprietary LLM baseline.
Machine Learning, ICML
1 Introduction
Accurate assessment of mental health risk is foundational to effective psychiatric and therapeutic care. Clinicians routinely integrate heterogeneous signals—sleep, activity, metabolic health, self-reported questionnaires, and lived context—into qualitative judgments about patient mood and risk. Automating or augmenting this process remains difficult, particularly when predictions must be discrete, interpretable, and temporally grounded.
Recent advances in large language models (LLMs) suggest promise in reasoning over structured and semi-structured health data. However, most prior work evaluates LLMs on open-ended clinical question answering rather than calibrated risk scoring. Furthermore, little work examines whether general-purpose LLMs can reliably predict longitudinal mood trajectories from compact behavioral feature sets.
This paper makes two primary contributions:
  • We introduce PEARL, a longitudinal behavioral healthcare dataset with monthly observations over four years, including physiological, behavioral, and self-reported mental health signals.
  • We propose HARBOR, a Behavioral Health–aware LLM trained to predict a clinically interpretable mood score (HRS) and demonstrate its superiority over classical models and proprietary LLMs.
Our goal is not to replace clinicians, but to explore whether structured, clinically grounded LLMs can serve as reliable decision-support tools in behavioral healthcare.
2 PEARL
PEARL is a small but deeply curated longitudinal dataset consisting of monthly observations from three adult patients over four years (48 months per patient, 144 total samples). Each data point consists of the following features:
  • Time and activity signals: sleep duration, step count, calories consumed and burned
  • Physiological markers: glucose, vitamin D, cholesterol, thyroid-stimulating hormone
  • Body composition: weight, body fat percentage
  • Behavioral proxies: number of photos taken, location entropy
  • Financial context: monthly expenses normalized by income
  • Clinical questionnaires: PHQ-9 and GAD-7
Each sample is paired with a self-evaluated mood score on a Likert scale from 3-3 (severe depression) to +3+3 (mania), which we refer to as the Harbor Risk Score (HRS).
Table 1: Main Results under Default Evaluation Settings
Method Accuracy Macro F1 Pearson Corr. Spearman Corr.
LogReg (L1) 0.50 0.30 0.82 0.83
LogReg (L2) 0.54 0.33 0.85 0.85
Random Forest 0.54 0.33 0.85 0.85
GPT-5.2 0.23 0.19 0.79 0.81
Claude 4.5 Sonnet 0.27 0.20 0.32 0.42
Grok 4.1 0.27 0.17 0.79 0.80
Gemini 3 Pro 0.29 0.26 0.80 0.83
HARBOR (Ours) 0.69 0.63 0.91 0.91
2.1 Ethical Considerations
Patients differ in ethnicity, gender, and socioeconomic background. No identifying information is included. All data was collected with informed consent and anonymized prior to use.
2.2 Dataset Splits
Unless otherwise stated, the default split consists of 48 training, 48 validation, and 48 test samples. We also evaluate alternative splits by patient identity and temporal ordering as part of our ablation studies.
3 HARBOR
HARBOR is initialized from a 20B-parameter open-source GPT-style checkpoint. The model is adapted to behavioral healthcare through a three-stage process: mid-training, supervised fine-tuning, and reinforcement learning.
3.1 Mid-Training
We perform mid-training on a curated corpus of psychiatry, psychology, and therapy textbooks, along with non-fiction behavioral health literature. This stage focuses on domain adaptation while preserving general language capabilities.
3.2 Fine-Tuning
Supervised Fine-Tuning (SFT). We generate structured question–answer pairs from domain textbooks and clinical guidelines, focusing on symptom interpretation, mood classification, and longitudinal reasoning.
Reinforcement Learning (RL). We apply reinforcement learning to encourage consistency, calibration, and adherence to the HRS scale. Rewards emphasize agreement with expert-aligned reasoning and penalize extreme or inconsistent predictions.
3.3 Self-Taught Reasoning
To improve structured reasoning over tabular inputs, we employ a self-taught reasoning (STaR) approach, where the model iteratively generates and refines its own reasoning traces during training.
4 Experiments and Results
4.1 Baselines
We compare HARBOR against:
  • Logistic Regression with L1 and L2 regularization
  • Random Forest
  • Proprietary LLMs: GPT-5.2, Claude 4.5 Sonnet, Grok 4.1, and Gemini 3 Pro
4.2 Evaluation Metrics
We report Accuracy, Macro-F1, Pearson correlation, and Spearman rank correlation between predicted and ground-truth HRS.
Default Prompt Used for Language Model Evaluation You are Harbor: Holistic Adaptive Risk assessment model for BehaviORal healthcare, a clinical decision-support assistant. Your task is to estimate a single discrete mood / risk score on an integer scale from -3 to +3 based on behavioral, physiological, and self-reported features. The scale is defined as follows: -3 = severe depression and unable to function or work, -2 = moderate depression with significant impairment, -1 = mild depressive symptoms, 0 = neutral or stable mood, +1 = mildly elevated mood, +2 = moderate mania or hypomania with impaired judgment or functioning, +3 = severe mania and unable to function or work. You will receive one independent example with comma-separated features in this exact order: time, sleep_minutes, calories_intake_kcal, calories_burned_kcal, num_steps, labs_glucose, labs_vitd, labs_cholestrol, labs_tsh, weight, body_fat_percent, num_pictures_taken, location, monthly_expense_by_income, phq_9, gad_7. All values are factual observations; phq_9 and gad_7 are validated clinical screening scores. Using clinical reasoning and weighing sleep, activity, metabolic health, anxiety/depression scores, and behavioral signals, infer the most likely overall mood state; do not assume any time-series context and treat the example independently. Output rules: return exactly one integer between -3 and +3 (inclusive), with no explanation, no extra text, and no formatting---only the number.
Figure 1: Prompt used for default evaluation of language models, including HARBOR and proprietary LLM baselines. Unicode minus signs are avoided for LaTeX compatibility.
Table 2: Accuracy under Different Prediction Horizons
Model $t_{0}$ (Current) $t_{-1}$ (1 Month) $t_{-3}$ (3 Months)
LogReg (L1) 0.50 0.43 0.35
LogReg (L2) 0.54 0.46 0.38
Random Forest 0.54 0.47 0.40
GPT-5.2 0.23 0.21 0.18
Claude 4.5 Sonnet 0.27 0.24 0.20
Grok 4.1 0.27 0.25 0.21
Gemini 3 Pro 0.29 0.26 0.23
HARBOR (Ours) 0.69 0.61 0.52
Table 3: Accuracy vs. Number of In-Context Examples (LLMs Only)
Model 0-shot 6-shot 48-shot
GPT-5.2 0.23 0.26 0.30
Claude 4.5 Sonnet 0.27 0.30 0.34
Grok 4.1 0.27 0.29 0.33
Gemini 3 Pro 0.29 0.33 0.37
HARBOR (Ours) 0.69 0.70 0.72
Table 4: Accuracy under Different Inference Modes (LLMs Only)
Model All at Once One by One
GPT-5.2 0.23 0.26
Claude 4.5 Sonnet 0.27 0.29
Grok 4.1 0.27 0.30
Gemini 3 Pro 0.29 0.32
HARBOR (Ours) 0.69 0.70
Table 5: Accuracy under Different Aggregation Strategies
Model Single Prediction Avg (5) Majority Vote (5)
GPT-5.2 0.23 0.25 0.26
Claude 4.5 Sonnet 0.27 0.29 0.30
Grok 4.1 0.27 0.30 0.31
Gemini 3 Pro 0.29 0.32 0.33
HARBOR (Ours) 0.69 0.71 0.72
Table 6: Accuracy under Different Dataset Split Strategies
Model Random Split Time-based Split Patient-based Split
LogReg (L2) 0.54 0.45 0.41
Random Forest 0.54 0.46 0.42
GPT-5.2 0.23 0.21 0.19
Claude 4.5 Sonnet 0.27 0.24 0.22
Grok 4.1 0.27 0.25 0.23
Gemini 3 Pro 0.29 0.27 0.25
HARBOR (Ours) 0.69 0.60 0.56
4.3 Default Evaluation Setting
Unless otherwise stated, all results are reported under a common default evaluation setting. Models are trained to predict the current-month Harbor Risk Score ($t_{0}$) using the full feature set described in Section 2. The dataset is split randomly into 48 training, 48 validation, and 48 test samples.
For language models, predictions are generated in a single batch over the entire test set using zero-shot prompting. A single prediction is produced per instance without aggregation or voting. Traditional machine learning baselines are trained using the training split with hyperparameters selected on the validation set and evaluated once on the held-out test set.
This default configuration is used for the main comparison across all methods. Variations along prediction horizon, prompting strategy, inference procedure, aggregation method, and dataset split are explored in the ablation studies.
4.4 Results
Table 1 summarizes performance under the default evaluation setting. Traditional machine learning models outperform off-the-shelf proprietary LLMs, suggesting that generic language models struggle to produce calibrated discrete risk scores from compact structured inputs. Among these baselines, logistic regression achieves the strongest performance, reflecting the small-data regime and the relatively linear relationship between features and mood labels. In contrast, HARBOR substantially outperforms all baselines across all metrics, achieving a 15-point absolute improvement in accuracy over the best traditional model. Notably, HARBOR also exhibits higher Pearson and Spearman correlations, indicating improved ordinal consistency and temporal calibration rather than simply better pointwise classification.
5 Ablation Studies
We evaluate five ablation dimensions: prediction horizon, number of in-context examples, inference mode, aggregation strategy, and dataset split strategy. For brevity and clarity, we report accuracy only in this section. Unless otherwise stated, all other experimental settings follow the default configuration described in Section 4.3. Full results, including additional metrics, will be released alongside the PEARL dataset.
5.1 Prediction Horizon
We first study the effect of prediction horizon by evaluating models on current-month mood prediction ($t_{0}$), next-month prediction ($t_{-1}$), and three-month-ahead prediction ($t_{-3}$). Accuracy degrades across all methods as the prediction horizon increases, reflecting the inherent uncertainty of long-term mood forecasting. Traditional models show sharp performance drops beyond the current month. HARBOR remains substantially more robust, retaining meaningful predictive signal even at a three-month horizon, suggesting improved temporal abstraction rather than simple pattern matching.
5.2 Number of In-Context Examples
We evaluate the impact of few-shot prompting on LLM performance by varying the number of in-context examples. Few-shot prompting improves all LLM baselines, but gains are modest and saturate quickly. Even with full training-set context, proprietary LLMs fail to approach traditional baselines. HARBOR benefits marginally from additional examples, indicating that most task-relevant structure is already internalized during training rather than inferred at inference time.
5.3 Inference Mode
We compare batch inference (all test samples predicted in a single prompt) with independent per-sample inference. Independent inference consistently improves accuracy for LLMs, suggesting that batch prompts may introduce cross-example interference. HARBOR shows minimal sensitivity to inference mode, indicating stronger per-sample calibration and reduced reliance on prompt context.
5.4 Aggregation Strategy
We examine whether aggregating multiple stochastic predictions improves robustness. Aggregation provides modest but consistent gains, particularly for LLMs with higher output variance. Majority voting slightly outperforms averaging, indicating discrete-mode stability. HARBOR benefits less from aggregation, reflecting more deterministic and stable predictions.
5.5 Dataset Split Strategy
Finally, we evaluate robustness to different dataset partitioning strategies. Performance drops under time-based and patient-based splits across all models, highlighting the difficulty of generalization in behavioral health. However, HARBOR exhibits significantly smaller degradation, suggesting improved robustness to distributional shift across both time and individuals.
6 Analysis
Several trends emerge from our experiments. First, off-the-shelf LLMs perform poorly despite strong general reasoning capabilities, suggesting that structured clinical risk scoring requires domain-specific adaptation. Second, traditional models benefit from the small dataset regime but plateau due to limited representational capacity. HARBOR combines domain knowledge with structured reasoning, enabling more calibrated and temporally consistent predictions.
We also observe that HARBOR degrades more gracefully under temporal and patient-based splits, indicating improved generalization across time and individuals.
7 Related Work
Risk Assessment in Psychiatry.
The challenge of predicting mental health outcomes has long been recognized in psychiatry. Classical work by Meehl demonstrated that simple statistical models can outperform clinical judgment in behavioral prediction, a result that has replicated across decades of clinical domains (meehl1954clinical). More recently, large-scale meta-analyses have shown that traditional psychiatric risk factors—particularly for suicide—have limited predictive power, motivating the use of machine learning–based risk models (franklin2017risk). In contrast to unstructured clinical judgment, structured risk scores such as the National Early Warning Score 2 (NEWS2) have seen widespread adoption in general medicine by aggregating physiological signals into an interpretable, discrete score for clinical decision-making (royal2017news2). However, comparable standardized scoring systems for behavioral health remain limited.
Structured and Longitudinal Behavioral Health Data.
Recent work has explored predictive modeling using structured electronic health records (EHRs), demonstrating improved performance for outcomes such as psychiatric readmission and suicide attempts (simon2018suicide; kessler2020predicting). Beyond EHRs, advances in mobile sensing and digital phenotyping have enabled continuous, longitudinal measurement of behavioral signals such as sleep, activity, mobility, and self-reported mood (onella2016mobile). Publicly released datasets capturing such signals have supported modeling of mood dynamics and relapse risk in real-world settings (pratap2020all; torous2021digital). These studies highlight the importance of combining physiological, behavioral, and self-report features—an approach directly reflected in the PEARL dataset.
Language Models in Clinical Decision Support.
Large language models (LLMs) have recently been explored for clinical applications, including medical question answering, summarization, and decision support (singhal2023large). Early studies suggest that LLMs can perform competitively on medical reasoning benchmarks, yet their reliability for calibrated risk prediction remains unclear (nori2023capabilities). In psychiatry, LLMs have been proposed for tasks such as mental health screening, therapy assistance, and patient engagement, but existing evaluations remain limited in scale and standardization (bickmore2018automated; torous2023trustworthy). Importantly, prior work largely focuses on free-text interaction rather than discrete, interpretable risk scoring.
Positioning of This Work.
HARBOR builds on these lines of research by combining structured behavioral health features with domain-adapted language modeling to produce a calibrated, discrete mood risk score. Unlike prior LLM-based systems, HARBOR is explicitly evaluated against traditional machine learning baselines and standardized across multiple ablation dimensions, positioning it as a step toward interpretable and reproducible behavioral health risk assessment.
8 Conclusion and Future Work
We introduce HARBOR, a Behavioral Health–aware LLM, and PEARL, a longitudinal behavioral healthcare dataset. Our results demonstrate that domain-adapted language models can significantly outperform both classical models and general-purpose LLMs in mood risk assessment.
Future work includes expanding PEARL to more patients, increasing temporal resolution to daily or hourly predictions, incorporating passive sensor data, and exploring HARBOR as a training and decision-support tool for clinicians and mental health professionals.